# Fair Document Valuation in LLM Summaries via Shapley Values

### v3

@@ 1 Introduction
Over the last few years, all major search engines have adopted these developments and integrated LLMs into their search infrastructures, e.g., Microsoft’s Bing AI (44) and Google’s AI Overview (22). Both these tools provide a summary response to a user’s search query with references or links to relevant websites for informational queries. [^1] Search queries are typically categorized into three groups – (1) informational, where the user is seeking information on a topic, (2) navigational, where the user is simply using the search engine to navigate to a website, and (3) transactional, where the user seeks to perform some transaction, e.g., a purchase. LLM-based summaries are mainly shown for informational queries. Further, new LLM-based search engines such as OpenAI’s ChatGPT Search (47) and Perplexity AI (50) have emerged and grown rapidly in the last few years. The impact of this transformation goes beyond web search and extends to Q&A and informational websites (e.g., Reddit) and e-commerce websites (e.g., Amazon, Best Buy). For example, Amazon now shows an LLM-based summary of all the reviews for a product on the main product page. Customers can further access the source reviews the summary is built on, if needed; see Figure 1 for an example. In addition to broad overviews of the reviews, consumers can query Amazon’s AI-based search tool Rufus for specific information related to the product, and it provides a response based on the information in the reviews/product page (59; 43). Examples of all these applications are illustrated in Web Appendix §A.
@@
However, the rise of LLM-based automated summarization tools has led to significant concerns among content creators and information providers. In a traditional search model, the source documents (and their corresponding producers) stand to gain traffic and reputation by being shown in response to search queries in organic listings. [^2] Both higher traffic and better reputation can usually be monetized through advertising and/or subscription revenues. In contrast, LLM-powered search delivers a synthesized answer directly in the search interface, allowing users to obtain information without visiting the source websites. For example, after the introduction of Google’s AI overview, website owners have expressed reservations over the loss of traffic and revenues (4; 63; 7). Similarly, on e-commerce and Q&A websites, individual users typically seek to develop a reputation for expertise and knowledge by providing meaningful answers or reviews and are often rewarded with badges, elite status, or free products in return. [^3] For example, sites like Reddit use a “Karma” system which captures how helpful/unhelpful users found a user’s responses (54). Similarly, Yelp recognizes high-quality reviewers by awarding them an “Elite” badge shown publicly on their profile, which in turn gives them access to attend exclusive events in the local community as well as other perks (6). However, if consumers stop reading individual reviews/answers and rating them, the reputation systems on these platforms would cease to function (56), and reviewers/content generators on these sites will lose the opportunity to monetize their content. Thus, the shift to LLM-based summaries threatens the revenue streams of content creators and raises concerns about the uncompensated use of their data. [^4] Indeed, several major publishers such as The New York Times have started to restrict or revoke AI access to their content, citing concerns that their contributions are being undervalued and misused without fair compensation (24). If content creators are not fairly compensated for their content, then they will have no incentive to provide content to platforms and search engines. This can lead to degradation in the quality of AI-powered summaries and failure of platforms’ reputation systems over time. Thus, while platforms may see short-term benefits from avoiding attribution or compensation to content creators and websites, this is not a sustainable equilibrium in the long term.
@@
We benchmark the proposed Cluster Shapley algorithm against two simple attribution rules-equal attribution and relevance-weighted attribution-and three widely used approximation methods: Monte Carlo (42), Truncated Monte Carlo (20), and Kernel SHAP (40). The results yield two main insights. (1) Off-the-shelf approximation algorithms are empirically far from optimal in the context of LLM-based summarization. These algorithms, originally designed for settings with independent contributors, fail to exploit the rich semantic structure of textual data and therefore deliver poor efficiency–accuracy trade-offs. In contrast, Cluster Shapley substantially improves the efficient frontier: for instance, our method requires only 20–40 unique permutations (out of 255 total) to reach an MAE of 0.20, whereas other approximation methods need more than 110 permutations to achieve comparable accuracy. This finding underscores the importance of developing Shapley approximation algorithms specifically tailored for LLM summarization-leveraging embedding-based similarity and other contextual structures-and points to a promising direction for future research. (2) Simple attribution rules are highly biased, raising fairness concerns in practical applications. Equal attribution, despite its negligible computational cost, results in an MAPE exceeding 150%, and even relevance-weighted attribution remains above 30%. These large deviations from the fair Shapley baseline demonstrate that simplistic heuristics, though computationally cheap, can substantially distort individual document contributions. Together, these findings highlight that more effective, structure-aware Shapley approximations are both feasible and necessary: they enable fairer value attribution with affordable computational cost.
@@ 2 Related Literature
Our research relates to and contributes to multiple streams of work, including game theory, LLMs, and business applications of AI systems. First, our work relates to the growing literature on LLM-based summarization. Recent research on summarization has shown that summaries produced by LLMs like GPT-4 achieve comparable or superior factual accuracy, coherence, and overall quality compared to human annotators in news summarization tasks (52). Indeed, text summarization has evolved from traditional extractive methods like TextRank (45) to more sophisticated abstractive approaches powered by LLMs. RAG frameworks (36), which combine information retrieval with LLM generation, have shown significant improvements in factual accuracy and information currency (19; 31). Recent developments include GraphRAG (14), which enhances retrieval performance using graph-based representations. While summarization technologies continue to evolve rapidly, our work addresses the fundamental challenge of document valuation that persists across summarization approaches. The document valuation framework we develop is designed to be agnostic to specific summarization techniques, ensuring its applicability even as LLM and RAG technologies advance.
@@
Next, our work builds on the game-theoretic concept of Shapley values (62), which has gained significant traction in machine learning for quantifying feature importance (40) and data valuation (30; 20). [^6] Apart from the Shapley value approach, alternative approaches for calculating the value of contributors in cooperative tasks have also been proposed. Some of the well-known approaches include the leave-one-out approach (10) and the influence function approach (3; 27; 25). However, these methods, neglecting high-order interactions, may result in undesirable data valuation. For example, when there are two or more similar contributors, and tend to assign nearly zero value to similar documents. For example, consider a scenario where we have three documents, A, B, and C, where A and B have the same high-quality content and C contributes nothing to the overall value. Intuitively, we should expect A and B to have the same valuation and C to have zero value. However, the leave-one-out approach will assign zero value to both A and B – if we leave A out and calculate the value generated by B and C, the value would almost be the same as that with A (and vice-versa for B). Thus, if we believe this approach and assume that both A and B have no incremental value and drop them both from the set of contributors, the overall value generated by the set will go to zero! In sum, these approaches are unable to correctly account for the marginal contributions of each player. This is the reason why Shapley has emerged as the dominant paradigm for valuing contributors in cooperative settings. While prior studies have applied Shapley values primarily to supervised learning tasks and feature attribution, our work represents, to our knowledge, the first application of Shapley values to document valuation in LLM-based summarization systems. Our work also relates to the work on approximation algorithms for computing Shapley values since it is well-known that computing exact Shapley values is computationally challenging. Without additional structural assumptions on the value function beyond bounded range, any unbiased estimator must rely on sampling permutations or coalitions; simple Monte Carlo (random-permutation) estimators are unbiased and achieve concentration rates that match information-theoretic lower bounds up to constants (5; 41). Specifically, with value range bounded by $[a,b]$, Hoeffding-style arguments yield a sample complexity of $O\!\left(\frac{(b-a)^{2}}{\varepsilon^{2}}\log\frac{1}{\delta}\right)$ to obtain an $\varepsilon$-accurate estimate with probability $1-\delta$, and Maleki et al. (41) give matching lower bounds. Alternative speedups such as linear regression surrogates, kernel/coalitional weighting, or grouping assumptions can reduce computation but introduce modeling assumptions that may bias attribution when violated (18; 2; 20).
@@ 3 Problem Definition
The quality or performance of a summarization is denoted by $v(q,A(q,S_{q}))$. Intuitively, this score captures the extent to which the user finds the summary useful or valuable. The performance score $v$ can be treated as a black-box that takes the query and summary as input and returns a score. In practice, $v(q,A(q,S_{q}))$ can be obtained in a multitude of ways. It could be actual scores collected from user surveys on how helpful they find a given summary to be (e.g., rating of helpfulness, fraction of upvotes). Alternatively, it could be helpfulness scores based on an LLM model, where an independent LLM agent does the scoring instead of human agents. This can be a viable option in settings where collecting user responses is costly and/or slow; indeed, recent research has shown that LLM ratings tend to align with user ratings in many situations (32; 8). It could also represent implicit helpfulness scores based on user behavior, which are commonly used in the information retrieval and search literature to measure the relevance of a given document/link, e.g., whether the user clicked on the summary, the time spent reading the summary (38; 71).
@@
    Summarization Procedure Agnostic: The approach should be agnostic to the specifics of the summarization process, $A(\cdot)$, used by the platform. That is, it should generalize across RAG systems. While we present a standard RAG implementation in $\S$6.1, numerous alternatives exist-from simpler methods to more sophisticated frameworks like TextRank, GraphRAG, and DRAG (45; 14; 74). As LLM technologies rapidly evolve, our document valuation framework is designed to remain effective, regardless of advancements in summarization, ensuring broad applicability across current and future implementations.
@@ 4.1 Shapley Value
A natural concern in attribution systems is whether document providers can strategically manipulate their content to inflate their assigned value. This issue is not new: even in traditional search engines like Google, publishers have long engaged in “mimicking the winner” strategies, which create content that resembles highly ranked pages to improve visibility. While such behavior is theoretically possible, it is difficult to execute in practice. Modern LLM-powered platforms (e.g., Perplexity, ChatGPT, Google’s AI Overview) do not disclose their specific RAG systems to external document providers. Moreover, these systems do not present explicit ranked lists but instead synthesize information into summaries, making it harder for publishers to infer their position or influence outcomes. Retrieval in such systems is governed by proprietary, black-box algorithms that incorporate a range of dynamic signals, including semantic embeddings, recency, diversity, and user engagement. This level of opacity is a deliberate design choice, observed not only in LLM systems but also in recommender platforms and digital advertising ecosystems, where hiding algorithmic details is a known strategy to deter gaming (35; 58).
@@ 4.3 Approximating Shapley Value: Cluster Shapley Approach
Text embedding techniques convert large chunks of text-such as sentences, paragraphs, or documents-into numerical vectors that capture semantic information. Earlier embedding methods, such as Word2Vec (46) and GloVe (49), are based on shallow neural networks and co-occurrence statistics, learning word-level embeddings by predicting surrounding context words or factorizing word co-occurrence matrices. These embeddings typically represent each word with a fixed vector, independent of context. In contrast, modern LLM-based embeddings, such as those produced by OpenAI’s latest text-embedding models, are generated using Transformer-based architectures and are pretrained on massive text corpora via next-token prediction objectives. These newer embeddings are contextualized-meaning the vector for a word or sentence depends on its surrounding context-and are typically high-dimensional (e.g., 3072 dimensions in OpenAI’s text-embedding-3-large model). Unlike generative LLMs designed for tasks like chat or text generation, embedding models are optimized to produce semantically meaningful representations suitable for a wide range of downstream tasks. LLM-based embedding vectors have been successfully applied to a wide variety of discriminative tasks, including text classification, document retrieval, sentiment analysis, and predicting the attractiveness of news headlines (48; 69). In our setting, we leverage these embeddings to cluster similar documents before computing Shapley values, allowing us to reduce redundant calculations.
@@ 4.4.1 Approximation Error of Cluster Shapley
There exists a constant $L>0$ such that for any two documents $i,j\in S_{q}$ and for any coalition $S\subseteq S_{q}$ not containing $i$ or $j$, the difference in their marginal contributions is bounded by the embedding distance:
$\Big|\big(v(S\cup\{i\})-v(S)\big)-\big(v(S\cup\{j\})-v(S)\big)\Big|\leq L\,d(e_{i},e_{j}).$ (9)

### v4

@@ 1 Introduction
Over the last few years, all major search engines have adopted these developments and integrated LLMs into their search infrastructures, e.g., Microsoft’s Bing AI (52) and Google’s AI Overview (28). Both these tools provide a summary response to a user’s search query with references or links to relevant websites for informational queries. [^1] Search queries are typically categorized into three groups – (1) informational, where the user is seeking information on a topic, (2) navigational, where the user is simply using the search engine to navigate to a website, and (3) transactional, where the user seeks to perform some transaction, e.g., a purchase. LLM-based summaries are mainly shown for informational queries. Further, new LLM-based search engines such as OpenAI’s ChatGPT Search (55) and Perplexity AI (58) have emerged and grown rapidly in the last few years. The impact of this transformation goes beyond web search and extends to Q&A and informational websites (e.g., Reddit) and e-commerce websites (e.g., Amazon, Best Buy). For example, Amazon now shows an LLM-based summary of all the reviews for a product on the main product page. Customers can further access the source reviews the summary is built on, if needed; see Figure 1 for an example. In addition to broad overviews of the reviews, consumers can query Amazon’s AI-based search tool Rufus for specific information related to the product, and it provides a response based on the information in the reviews/product page (67; 51). Examples of all these applications are illustrated in Web Appendix §A.
@@
However, the rise of LLM-based automated summarization tools has led to significant concerns among content creators and information providers. In a traditional search model, the source documents (and their corresponding producers) stand to gain traffic and reputation by being shown in response to search queries in organic listings. [^2] Both higher traffic and better reputation can usually be monetized through advertising and/or subscription revenues. In contrast, LLM-powered search delivers a synthesized answer directly in the search interface, allowing users to obtain information without visiting the source websites. For example, after the introduction of Google’s AI overview, website owners have expressed reservations over the loss of traffic and revenues (8; 71; 11). Similarly, on e-commerce and Q&A websites, individual users typically seek to develop a reputation for expertise and knowledge by providing meaningful answers or reviews and are often rewarded with badges, elite status, or free products in return. [^3] For example, sites like Reddit use a “Karma” system which captures how helpful/unhelpful users found a user’s responses (62). Similarly, Yelp recognizes high-quality reviewers by awarding them an “Elite” badge shown publicly on their profile, which in turn gives them access to attend exclusive events in the local community as well as other perks (10). However, if consumers stop reading individual reviews/answers and rating them, the reputation systems on these platforms would cease to function (64), and reviewers/content generators on these sites will lose the opportunity to monetize their content. Thus, the shift to LLM-based summaries threatens the revenue streams of content creators and raises concerns about the uncompensated use of their data. [^4] Indeed, several major publishers such as The New York Times have started to restrict or revoke AI access to their content, citing concerns that their contributions are being undervalued and misused without fair compensation (30). If content creators are not fairly compensated for their content, then they will have no incentive to provide content to platforms and search engines. This can lead to degradation in the quality of AI-powered summaries and failure of platforms’ reputation systems over time. Thus, while platforms may see short-term benefits from avoiding attribution or compensation to content creators and websites, this is not a sustainable equilibrium in the long term.
@@
We benchmark the proposed Cluster Shapley algorithm against two simple attribution rules-equal attribution [^6] A natural question arises: can we simply resort to simpler heuristic allocations, such as uniform distribution? Theoretical works have explored this possibility; for instance, Chen and Yin (12) demonstrates that under specific conditions, typically in large games with infinitesimal players, the Shapley value allocations can converge to uniform allocations. However, it does not hold for LLM summarization tasks. In our context, the “players” (source documents) are finite in number and highly heterogeneous in their semantic contribution. and relevance-weighted attribution-and three widely used approximation methods: Monte Carlo (49), Truncated Monte Carlo (26), and Kernel SHAP (47). The results yield two main insights. (1) Off-the-shelf approximation algorithms are empirically far from optimal in the context of LLM-based summarization. These algorithms, originally designed for settings with independent contributors, fail to exploit the rich semantic structure of textual data and therefore deliver poor efficiency–accuracy trade-offs. In contrast, Cluster Shapley substantially improves the efficient frontier: for instance, our method requires only 20–40 unique permutations (out of 255 total) to reach an MAE of 0.20, whereas other approximation methods need more than 110 permutations to achieve comparable accuracy. This finding underscores the importance of developing Shapley approximation algorithms specifically tailored for LLM summarization-leveraging embedding-based similarity and other contextual structures-and points to a promising direction for future research. (2) Simple attribution rules are highly biased, raising fairness concerns in practical applications. Equal attribution, despite its negligible computational cost, results in an MAPE exceeding 150%, and even relevance-weighted attribution remains above 30%. These large deviations from the fair Shapley baseline demonstrate that simplistic heuristics, though computationally cheap, can substantially distort individual document contributions. Together, these findings highlight that more effective, structure-aware Shapley approximations are both feasible and necessary: they enable fairer value attribution with affordable computational cost.
@@ 2 Related Literature
Our research relates to and contributes to multiple streams of work, including game theory, LLMs, and business applications of AI systems. First, our work relates to the growing literature on LLM-based summarization. Recent research on summarization has shown that summaries produced by LLMs like GPT-4 achieve comparable or superior factual accuracy, coherence, and overall quality compared to human annotators in news summarization tasks (60). Indeed, text summarization has evolved from traditional extractive methods like TextRank (53) to more sophisticated abstractive approaches powered by LLMs. RAG frameworks (43), which combine information retrieval with LLM generation, have shown significant improvements in factual accuracy and information currency (25; 38). Recent developments include GraphRAG (20), which enhances retrieval performance using graph-based representations. While summarization technologies continue to evolve rapidly, our work addresses the fundamental challenge of document valuation that persists across summarization approaches. The document valuation framework we develop is designed to be agnostic to specific summarization techniques, ensuring its applicability even as LLM and RAG technologies advance.
@@
Next, our work builds on the game-theoretic concept of Shapley values (70), which has gained significant traction in machine learning for quantifying feature importance (47) and data valuation (37; 26). [^7] Apart from the Shapley value approach, alternative approaches for calculating the value of contributors in cooperative tasks have also been proposed. Some of the well-known approaches include the leave-one-out approach (15) and the influence function approach (5; 33; 31). However, these methods, neglecting high-order interactions, may result in undesirable data valuation. For example, when there are two or more similar contributors, and tend to assign nearly zero value to similar documents. For example, consider a scenario where we have three documents, A, B, and C, where A and B have the same high-quality content and C contributes nothing to the overall value. Intuitively, we should expect A and B to have the same valuation and C to have zero value. However, the leave-one-out approach will assign zero value to both A and B – if we leave A out and calculate the value generated by B and C, the value would almost be the same as that with A (and vice-versa for B). Thus, if we believe this approach and assume that both A and B have no incremental value and drop them both from the set of contributors, the overall value generated by the set will go to zero! In sum, these approaches are unable to correctly account for the marginal contributions of each player. This is the reason why Shapley has emerged as the dominant paradigm for valuing contributors in cooperative settings. While prior studies have applied Shapley values primarily to supervised learning tasks and feature attribution, our work represents, to our knowledge, the first application of Shapley values to document valuation in LLM-based summarization systems. Our work also relates to the work on approximation algorithms for computing Shapley values since it is well-known that computing exact Shapley values is computationally challenging. Without additional structural assumptions on the value function beyond bounded range, any unbiased estimator must rely on sampling permutations or coalitions; simple Monte Carlo (random-permutation) estimators are unbiased and achieve concentration rates that match information-theoretic lower bounds up to constants (9; 48). Specifically, with value range bounded by $[a,b]$, Hoeffding-style arguments yield a sample complexity of $O\!\left(\frac{(b-a)^{2}}{\varepsilon^{2}}\log\frac{1}{\delta}\right)$ to obtain an $\varepsilon$-accurate estimate with probability $1-\delta$, and Maleki et al. (48) give matching lower bounds. Alternative speedups such as linear regression surrogates, kernel/coalitional weighting, or grouping assumptions can reduce computation but introduce modeling assumptions that may bias attribution when violated (24; 4; 26).
@@
Finally, our work relates to the literature on news aggregators in marketing and economics. News aggregators function similarly to AI-based summarization, since users may consume content directly on the aggregator’s platform without visiting the publishers’ websites. A stream of research has examined whether this substitutive effect reduces traffic to publishers or whether aggregators can also serve as a discovery channel that increases exposure. For instance, Mayzlin and Yoganarasimhan (50) find that blogs can build an audience by effectively promoting/linking to rival sites under certain conditions. Dellarocas et al. (19) develop a game-theoretic model and show that content aggregators can both benefit and harm content creators and consumers. Jeon and Nasr (36) investigate competition among online newspapers in the presence of aggregators. Empirically, both Calzada and Gil (7) and Athey et al. (3) find that news aggregators can exert a market-expansion effect, increasing visits to news outlets, especially for smaller publishers. Further, recent work by Amaldoss and Du (2) explores how publishers might collaborate and compete with news aggregators, and Song and Manchanda (72) empirically examine the effects of carrying news on user engagement with non-news content on social media.
This perspective also closely aligns with the emerging literature on the creator economy, which examines how platforms aggregate independent content providers through bundling and revenue-sharing mechanisms. For instance, Bhargava (6) develops a model of platform design, demonstrating that revenue sharing is not merely a zero-sum game between platforms and creators; rather, an optimal sharing rule effectively incentivizes ecosystem supply and enhances total welfare. Complementing this, recent studies on bundle pricing (84) suggest that aggregating high-variety goods can reduce demand volatility and improve monetization efficiency. Recent empirical work by Wu and Zhu (77) highlights the critical role of incentive design and demonstrates that performance-based revenue-sharing contracts, as opposed to fixed compensation, significantly boost novelty.
Given that AI-based summarization is functionally similar to content aggregators and bundling, our Cluster Shapley approach for valuing source documents can also be used by content aggregators and newspapers/content websites to formalize revenue-sharing arrangements.
@@ 3 Problem Definition
The quality or performance of a summarization is denoted by $v(q,A(q,S_{q}))$. Intuitively, this score captures the extent to which the user finds the summary useful or valuable. The performance score $v$ can be treated as a black-box that takes the query and summary as input and returns a score. In practice, $v(q,A(q,S_{q}))$ can be obtained in a multitude of ways. It could be actual scores collected from user surveys on how helpful they find a given summary to be (e.g., rating of helpfulness, fraction of upvotes). Alternatively, it could be helpfulness scores based on an LLM model, where an independent LLM agent does the scoring instead of human agents. This can be a viable option in settings where collecting user responses is costly and/or slow; indeed, recent research has shown that LLM ratings tend to align with user ratings in many situations (39; 13). It could also represent implicit helpfulness scores based on user behavior, which are commonly used in the information retrieval and search literature to measure the relevance of a given document/link, e.g., whether the user clicked on the summary, the time spent reading the summary (45; 81).
@@
    Summarization Procedure Agnostic: The approach should be agnostic to the specifics of the summarization process, $A(\cdot)$, used by the platform. That is, it should generalize across RAG systems. While we present a standard RAG implementation in $\S$6.1, numerous alternatives exist-from simpler methods to more sophisticated frameworks like TextRank, GraphRAG, and DRAG (53; 20; 85). As LLM technologies rapidly evolve, our document valuation framework is designed to remain effective, regardless of advancements in summarization, ensuring broad applicability across current and future implementations.
@@ 4.1 Shapley Value
A natural concern in attribution systems is whether document providers can strategically manipulate their content to inflate their assigned value. This issue is not new: even in traditional search engines like Google, publishers have long engaged in “mimicking the winner” strategies, which create content that resembles highly ranked pages to improve visibility. While such behavior is theoretically possible, it is difficult to execute in practice. Modern LLM-powered platforms (e.g., Perplexity, ChatGPT, Google’s AI Overview) do not disclose their specific RAG systems to external document providers. Moreover, these systems do not present explicit ranked lists but instead synthesize information into summaries, making it harder for publishers to infer their position or influence outcomes. Retrieval in such systems is governed by proprietary, black-box algorithms that incorporate a range of dynamic signals, including semantic embeddings, recency, diversity, and user engagement. This level of opacity is a deliberate design choice, observed not only in LLM systems but also in recommender platforms and digital advertising ecosystems, where hiding algorithmic details is a known strategy to deter gaming (42; 66).
@@ 4.3 Approximating Shapley Value: Cluster Shapley Approach
Text embedding techniques convert large chunks of text-such as sentences, paragraphs, or documents-into numerical vectors that capture semantic information. Earlier embedding methods, such as Word2Vec (54) and GloVe (57), are based on shallow neural networks and co-occurrence statistics, learning word-level embeddings by predicting surrounding context words or factorizing word co-occurrence matrices. These embeddings typically represent each word with a fixed vector, independent of context. In contrast, modern LLM-based embeddings, such as those produced by OpenAI’s latest text-embedding models, are generated using Transformer-based architectures and are pretrained on massive text corpora via next-token prediction objectives. These newer embeddings are contextualized-meaning the vector for a word or sentence depends on its surrounding context-and are typically high-dimensional (e.g., 3072 dimensions in OpenAI’s text-embedding-3-large model). Unlike generative LLMs designed for tasks like chat or text generation, embedding models are optimized to produce semantically meaningful representations suitable for a wide range of downstream tasks. LLM-based embedding vectors have been successfully applied to a wide variety of discriminative tasks, including text classification, document retrieval, sentiment analysis, and predicting the attractiveness of news headlines (56; 79). In our setting, we leverage these embeddings to cluster similar documents before computing Shapley values, allowing us to reduce redundant calculations.
@@ 4.4.1 Approximation Error of Cluster Shapley
There exists a constant $L>0$ such that for any two documents $i,j\in S_{q}$ satisfying the condition $d(e_{i},e_{j})\leq\epsilon$, the difference in their marginal contributions is bounded by their embedding distance:
$\Big|\big(v(S\cup\{i\})-v(S)\big)-\big(v(S\cup\{j\})-v(S)\big)\Big|\leq L\,d(e_{i},e_{j}),$ (9)
for any coalition $S\subseteq S_{q}\setminus\{i,j\}$. Note that since $d(e_{i},e_{j})\leq\epsilon$, this specifically implies the difference is bounded by $L\epsilon$ within the cluster.
Under Assumption 1, if two documents lie in the same cluster ($d(e_{i},e_{j})\leq\epsilon$), their marginal contributions to any coalition differ by at most $L\epsilon$. Intuitively, this assumption states that semantically similar documents should exhibit similar marginal contributions across all possible coalitions. The assumption is both theoretically reasonable and practically essential: documents with similar content naturally provide comparable value when added to any subset of documents, ensuring that our clustering approach meaningfully groups functionally equivalent items.